OpenStudy (anonymous):

What value of h makes the statement true? (17 + h) + (–9) = 17 + (5 + (–9)) A. –21 B. –5 C. 5 D. 21

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

not quite (17 + h) + (–9) = 17 + (5 + (–9)) 17 - 9 + h = 17 + 5 - 9 8 + h = 13 -8 -8 h = ?

OpenStudy (anonymous):

A?

OpenStudy (sleepyhead314):

13 - 8 = ?

OpenStudy (anonymous):

5